OpenStudy (anonymous):

sqrt -6 times sqrt -18 ?

OpenStudy (anonymous):

Do you mean, \[(\sqrt{-6})(\sqrt{-18})\]

OpenStudy (anonymous):

Are you familiar with complex numbers? \[(√-6)(√-18) = (6i)(18i)\] \[= 108i^2 - 108(-1) = -108 \]

OpenStudy (anonymous):

\[(i \sqrt{6})(i \sqrt{18})\] \[ -\sqrt{6}\sqrt{18}\] \[-\sqrt{108}\] \[-6\sqrt{3}\]

OpenStudy (anonymous):

\[ \sqrt{-6} \times \sqrt{-18}\]

OpenStudy (theviper):

Yes \[\LARGE{\sqrt{-6}\times\sqrt{-18}}\]\[\LARGE{=\sqrt{-6\times-18}}\]\[\LARGE{=\sqrt{108}}\]

OpenStudy (theviper):

@jricafort here are two \(i\) therefore it will become +

OpenStudy (zarkon):

jricafort is correct

OpenStudy (theviper):

why ?

OpenStudy (zarkon):

you are using a rule incorrectly

OpenStudy (anonymous):

@TheViper \[i ^{2} = -1, i ^{3} = -i, i ^{4} = 1, \]

OpenStudy (zarkon):

if \(a,b\ge 0\) then \[\sqrt{a}\sqrt{b}=\sqrt{ab}\] this is not true is a,b<0

OpenStudy (zarkon):

this is not true if a,b<0

OpenStudy (theviper):

\[\LARGE{\sqrt{-6}\times\sqrt{-18}}\]\[\LARGE{=\sqrt{-1}\times\sqrt{-1}\times\sqrt{6}\times\sqrt{18}}\]\[\LARGE{=(\sqrt{-1})^2}\times{108}\]\[\LARGE{=-1\times108=-108}\]

OpenStudy (theviper):

Sorry!

OpenStudy (anonymous):

thank you everyone(:

OpenStudy (anonymous):

No problem!

OpenStudy (theviper):

right @mayankdevnani :)

OpenStudy (mayankdevnani):

right @TheViper